Write five pairs of prime numbers less than 20 whose sum is divisible by 5.
(Hint: 3 + 7 = 10)

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Required pairs of prime numbers less than 20 are: (i) 2 + 3 = 5 (ii) 2 + 13 = 15 (iii) 11 + 9 = 20 (iv) 17 + 3 = 20 (v) 7 + 13 = 20